Prove that the series ∞n=1 n^n/n! diverges


1
Expert's answer
2021-02-24T06:39:43-0500

By the Inequality of arithmetic and geometric means we have:

"\sqrt{1\cdot (n-1)}\leq n/2" or "1\cdot (n-1)\leq (n/2)^2"

"\sqrt{2\cdot (n-2)}\leq n/2" , or "2\cdot (n-2)\leq (n/2)^2"

...

if n is even, the last inequality will be "n/2\leq n/2",

if n is odd, the last inequality will be "(n-1)/2\cdot (n+1)/2\leq (n/2)^2"

Multiplying all this inequalities (without the square roots), we obtain:

"(n-1)!\leq (n/2)^{n-1}" and

"n!\leq 2(n/2)^n"

Therefore, "\frac{n^n}{n!}\geq\frac{n^n}{2(n/2)^n}=2^{n-1}"

The values "2^{n-1}" are unbounded from above, therefore the sequence "{n^n}\/n!" is unbounded too and hence, it is divergent.